# Neural Transformer Backflow for Solving Momentum-Resolved Ground States of Strongly Correlated Materials

### v1

@@ Abstract
Strongly correlated materials, such as twisted transition-metal dichalcogenide homobilayers, host a variety of exotic quantum phases but remain notoriously difficult to solve due to strong interactions. We introduce a powerful neural network ansatz, Neural Transformer Backflow (NTB), formulated within a multi-band projection framework. It naturally enforces momentum conservation and enables efficient calculations of momentum-resolved ground states. NTB attains high accuracy on small systems and scales to higher bands and larger system sizes far beyond the reach of exact diagonalization. By evaluating observables such as the structure factor and momentum distribution, we show that NTB captures diverse correlated states in tMoTe2, including charge density waves, fractional Chern insulators, and anomalous Hall Fermi liquids, within a unified framework. Our approach paves the way for understanding and discovering novel phases of matter in strongly correlated materials.
@@
Introduction—. Recent advances in strongly correlated materials has lead to numerous experimental discoveof unconventional physics, such as Fractional Quantumries of exotic physics, such as Fractional Quantum Hall(FQH) states [1, 2], superconductivity [3, 4], and generalized Wigner Crystals [5, 6]. In particular, Moiré superlattices in twisted transition-metal dichalcogenide homobilayers (tTMDs) generate strong electron correlations from flat band structures [7], leading to novel states with nontrivial topology [8, 9, 10, 11]. External tuning parameters can drastically alter their behavior [12, 13], establishing tTMDs as a versatile platform for exploring correlated and topological quantum phenomena [14, 15].
@@
Meanwhile, recent advances in neural networks (NN) have demonstrated an alternative avenue to address such problem. With its remarkable representational power across all domains [23, 24, 25, 26], NN has emerged as a compelling candidate for solving many-body problems. A natural framework for applying NN in this context is the neural quantum state (NQS)  [27, 26, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. For fermionic systems, Neural Network Backflow(NNBF) has been proposed to encode fermionic statistics and strong correlation [47], inspiring later advancement for solving fermionic model Hamiltonians [26, 48, 49] and quantum chemistry [50, 51, 52, 53, 54].
While it has been shown that NQS can accurately capture many-body physics under the real-space formalism  [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70], incorporating NQS into the MBP formalism remains unexplored. In particular, despite the full Hilbert space without band truncation can be access in real space, enforcing momentum conservation becomes computationally expensive due to the need for symmetrization over all lattice translations. This makes the calculation of the energy spectrum of momentum resolved ground states, which contains important features such as degeneracies and energy gaps, computationally expensive. In contrast, in the MBP formalism, distinct momentum sectors are naturally decoupled due to momentum conservation.
In this work, we develop a powerful NQS ansatz, namely Neural Transformer Backflow(NTB), to solve the MBP Hamiltonian. Incorporating the high representation power of Transformer in NNBF, NTB extends the HF formalism to a many-body level, exceeding the performance of existing architectures. By employing NTB to solve the MBP Hamiltonian for moiré materials, we demonstrate that NTB not only agrees with ED results for small systems with a few bands, but also extends beyond the limitations of ED, enabling access to higher bands and larger system sizes. Furthermore, we also show that NTB can accurately capture intricate quantum phases over various of filling factors, including Charge Density Waves(CDW), Fractional Chern Insulator(FCI) and Anomalous Hall Fermi Liquids (AHFL), in moiré materials across a wide range of parameter regimes. This offers a unified approach for the simulation of strongly correlated materials within the MBP framework.
@@
The representability of NTB stems from three important designs. First, ${\bf h}_{n}$ can be regarded as a many-body bitstring enriched by the representational power of the Transformer. If ${\bf h}_{n}=x_{n}$ without backflow transformation and $N_{\rm hid}=1$, ${W}_{n}$ corresponds exactly to the mean field orbitals, reducing the architecture to the variational HF method. Thus, NTB can be viewed as a natural extension of Hartree–Fock, augmented by the expressive capacity of Transformer for strong correlation. Second, as the linear layers are designed per orbital, the architecture avoids a fully dense connection between the Transformer output and neural orbitals, in contrast to Ref. [49]. This design substantially decrease the total number of network parameter, which allows us to include more determinants to further enhancing the representability of the network within the computational capacity. Third, by taking the configuration-dependent Slater determinants on the neural orbitals via backflow transformation, the fermionic statistic is preserved, which further strengthens the representability. For the rest of this paper, we use a hidden dimension $N_{\rm hid}=32$, number of layers $N_{\rm layer}=2$ and number of attention heads $N_{\rm head}=16$.
@@
Moreover, we note that the inference time with momentum conservation scales with $\mathcal{O}(N_{e}^{3})$ for our method, as momentum is naturally conserved during sampling. This outperform the $\mathcal{O}(N_{e}^{4})$ scaling for real space approaches, where the the extra factor arises from averaging over lattice translations. We also note that transfer learning can be naturally implemented across different momentum sectors, which significantly reduces the optimization time.
@@
Discussions—.In this work, we have purposed a new network architecture NTB to solve the MBP Hamiltonians of strongly correlated materials. By calculating momentum-resolved ground states of fractional moiré materials, NTB has demonstrated excelling scalings among several architecture candidates. Moreover, it also goes beyond the limit of ED by including higher bands and larger system sizes, achieving a much lower ground state energy. By calculating observables such as $S({\bf{q}})$ and $n({\bf{k}})$, we demonstrate that NTB can capture a broad range of nontrivial topological and correlated quantum states. By enforcing momentum conservation during sampling, our method allows efficient calculation of momentum-resolved ground states across momentum sectors, providing a more efficient tool towards a more comprehensive understanding of strongly correlated materials. Future directions include exploring the calculation of topological quantities, such as the Chern number, within the band projected framework. Leveraging NTB’s scalability and the ability to probe a wide range of observables, our approach offers the potential to map out a global phase diagram for various quantum materials across different parameter regimes.

### v3

@@ Abstract
Strongly correlated materials host a rich variety of exotic quantum phases but remain challenging to solve due to strong interactions. We introduce the Neural Transformer Backflow (NTB) framework, a powerful neural-network ansatz formulated within a multi-band projection formalism. NTB is mean-field transcendental, parameter-efficient and fermionic intrinsic, exhibiting superior performance compared with existing neural ansatzes. By naturally enforcing momentum conservation, NTB enables direct computation of momentum-resolved many-body ground states, providing detailed access to degeneracies and energy gaps. It achieves high accuracy on small systems and scales efficiently to larger sizes and higher-band truncations far beyond the reach of exact diagonalization. We demonstrate the power of NTB in capturing diverse correlated phases in twisted MoTe2, including charge density waves, fractional Chern insulators, and anomalous Hall Fermi liquids, within a unified framework. This approach offers a generic, scalable route towards understanding and discovering quantum phases in strongly correlated materials.
@@
Introduction—. Recent advances in strongly correlated materials has lead to numerous experimental discoveries of unconventional physics, such as fractional quantum Hall (FQH) states [1, 2], superconductivity [3, 4], and generalized Wigner crystals [5, 6]. In particular, moiré superlattices in twisted transition-metal dichalcogenide homobilayers (tTMDs) generate strong electron correlations from flat band structures [7], leading to novel states with nontrivial topology [8, 9, 10, 11]. External tuning of parameters can drastically alter their behavior [12, 13], establishing tTMDs as a versatile platform for exploring correlated and topological quantum phenomena [14, 15].
@@
Meanwhile, recent advances in neural networks (NN) have demonstrated an alternative avenue to address such problem. With its remarkable representational power across all domains [23, 24, 25, 26], NN has emerged as a compelling candidate for solving many-body problems. A natural framework for applying NN in this context is the neural quantum state (NQS)  [27, 26, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. For fermionic systems, neural network backflow (NNBF) has been proposed to encode fermionic statistics and strong correlation [48], inspiring later advancement for solving fermionic model Hamiltonians [26, 49, 50] and quantum chemistry [51, 52, 53, 54, 55].
While it has been shown that NQS can accurately capture many-body physics under the real-space formalism  [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72], incorporating NQS into the MBP formalism remains unexplored. In particular, despite the full Hilbert space without band truncation can be access in real space, enforcing momentum conservation becomes computationally expensive due to the need for symmetrization over all lattice translations. This makes the calculation of the energy spectrum of momentum resolved ground states, which contains important features such as degeneracies and energy gaps, computationally expensive. In contrast, in the MBP formalism, distinct momentum sectors are naturally decoupled due to momentum conservation.
In this work, we develop a neural Transformer backflow (NTB) framework, to solve the MBP Hamiltonian. Our architecture is mean-field transcendental, parameter-efficient, and fermionic intrinsic, exhibits superior performance compared to existing neural ansatzes. We also introduce a new sampling technique, which allows the calculation of the momentum-resolved many-body ground-state spectrum, revealing important information on energy degeneracy and gaps. By employing NTB to solve the MBP Hamiltonian for moiré materials, we demonstrate that NTB not only agrees with exact diagonalization (ED) results for small systems with a few bands, but also extends beyond the limitations of ED, enabling access to higher bands and larger system sizes. Furthermore, we also show that NTB can accurately capture intricate quantum phases over various of filling factors, including charge density waves (CDW), fractional Chern insulator (FCI) and anomalous Hall Fermi liquids (AHFL), in moiré materials across a wide range of parameter regimes. This offers a unified approach for the simulation of strongly correlated materials within the MBP framework.
@@
The advancement and novelty in NTB architecture stem from three important properties:
  1. 1.
    Mean-field transcendental. By simply setting $N_{\rm hid}=N_{\rm det}=1$ and ${\bf h}_{n}=x_{n}$, ${W}_{n}$ corresponds exactly to the mean field orbitals, reducing the architecture to the same mathematical structure as HF method. This allows NTB to naturally encodes many-body effects into the mean-field framework. Powered by Transformer’s expressivity, NTB is able to capture many-body correlations, transcending the performance of mean-field methods.
  2. 2.
    Parameter-efficient. The linear layers are designed per orbital, which avoids a fully dense connection between the Transformer output and neural orbitals. This substantially decrease the total number of network parameter, allowing further augmentation of network performance by increasing $N_{\rm det}$.
  3. 3.
    Fermionic intrinsic. The final amplitude is obtained by taking configuration-dependent Slater determinants, which preserved fermionic statistics intrinsically, different from the previous work on direct application of Transformer.
For the rest of this paper, we use a hidden dimension $N_{\rm hid}=32$, number of layers $N_{\rm layer}=2$ and number of attention heads $N_{\rm head}=16$.
@@
Moreover, we note that the inference time with momentum conservation scales with $\mathcal{O}(N_{e}^{3})$ for our method, as momentum is naturally conserved during sampling. This is in contrast with the $\mathcal{O}(N_{e}^{4})$ scaling for wavefunction symmetrization approaches [78], where the the extra factor arises from averaging over lattice translations. We also note that transfer learning can be naturally implemented across different momentum sectors, which significantly reduces the optimization time.
@@
Discussions—. In this work, we have introduced a new neural-network backflow framework to solve the MBP Hamiltonians of strongly correlated materials. NTB is mean-field transcendental, parameter-efficient, and fermionic intrinsic, demonstrating excelling performances compared to existing architectures. Moreover, it goes beyond the limit of ED by including higher bands and larger system sizes, achieving a much lower ground state energy. By enforcing momentum conservation during sampling, our method allows efficient calculation of momentum-resolved ground states across momentum sectors, providing a more efficient tool towards a more comprehensive understanding of strongly correlated materials. In addition, by calculating observables such as $S({\bf{q}})$ and $n({\bf{k}})$, we demonstrate that NTB can capture a wide range of nontrivial topological and correlated quantum states. Future directions include exploring the calculation of superconductivity and topological quantities, such as the Chern number, within the band projected framework. Leveraging NTB’s scalability and the ability to probe a wide range of observables, our approach offers the potential to map out a global phase diagram for various quantum materials across different parameter regimes.